Circular No. 7960

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2002 Q2 (LINEAR) AND C/2002 Q3 (LINEAR) S. Adams, Lincoln Laboratory, Massachusetts Institute of Technology,

reports the discovery of two comets, both with coma diameter ∼ 90′′ and both in the same field and moving in the same direction on LINEAR images taken on Aug. 26.1 UT. Following posting on the NEO Confirmation Page, additional CCD observers have confirmed both objects to be cometary in appearance. On Aug. 26.9, P. Kušnirak (0.65-m f/3.6 reflector, Ondřejov) found C/2002 Q2 to be well condensed with a coma of diameter ∼ 20′′ and V = 17.7, whereas C/2002 Q3 appeared very diffuse without condensation and coma diameter ∼ 25′′. A. Galád (0.6-m reflector, Modra) reported C/2002 Q2 to be diffuse with m_{1} = 17.2, and S. P. Laurie (0.35-m reflector, Church Stretton, U.K.) reported the same object as diffuse with an 8′′ coma. P. R. Holvorcem and M. Schwartz (0.36-m reflector, Cottage Grove, OR) found C/2002 Q2 to be diffuse with a coma of diameter ∼ 17′′ and m_{1} = 17.1 on Aug. 27.3, whereas they noted C/2002 Q3 to be 2–2.5 mag fainter with a hint of diffuseness. The available astrometry, together with the preliminary parabolic orbital elements given below (by B. G. Marsden) that confirm the apparent relationship of the two comets (with C/2002 Q2 as the primary component), appear on MPEC 2002-Q37. The discovery observation and orbital elements for C/2002 Q2:

2002 UT  α_{2000}  δ_{2000}  m_{1}
Aug. 26.14383  17^{h}16^{m}24.^{s}15  +76^{o}49′54′′.5  18.8
T = 2002 Aug. 19.883 TT  ω = 127.^{o}932
Ω = 154.812 }2000.0
q = 1.29526 AU  i = 96.235

The discovery observation and orbital elements for C/2002 Q3:

2002 UT  α_{2000}  δ_{2000}  m_{1}
Aug. 26.14383  17^{h}04^{m}15.^{s}38  +76^{o}58′42′′.3  19.3
T = 2002 Aug. 20.612 TT  ω = 127.^{o}733
Ω = 154.696 }2000.0
q = 1.29744 AU  i = 96.366
2002 August 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green